You and your partner have become very interested in cross-country motorcycle racing and wish to purchase entry-level equipment. You have identified two alternative sets of equipment and gear. Package K has a first cost of $200,000, an operating cost of $7,500 per quarter, and a salvage value of $30,000 after its 2-year life. Package L has a first cost of $290,000 with a lower operating cost of $3,700 per quarter and an estimated $25,000 salvage value after its 4-year life. Which package offers the lower present worth analysis at an interest rate of 8% per year, compounded quarterly? The present worth of package K is $ and that of package L is $ .
Answer #1

Interest rate per quarter = 8/4 = 2%

n = 2years * 4 = 8 quarters

PW of project K = - 200,000 - 7500*(P/A, 2%, 8) + 30,000(P/F, 2%, 8)

= -200,000 - 7500*7.325 + 30,000* 0.8535

= - 229,332.5

Similarly,

For project L

n = 4 years *4 = 16 quarters

PW of project L = - 290,000 - 3700*(P/A, 2%, 16) + 25,000(P/F, 2%, 16)

= -290,000 - 3700*13.578 + 25,000* 0.7284

= - 322,028.6

As the PW are in negative numbers, so the package L offers the lower present worth analysis at an interest rate of 8% per year, compounded quarterly.